# What is the Plausible Memory Representation?

### v1

@@ Abstract
Recent work in computational psycholinguistics has revealed intriguing parallels between attention mechanisms and human memory retrieval, focusing primarily on Transformer architectures that operate on token-level representations. However, computational psycholinguistic research has also established that syntactic structures provide compelling explanations for human sentence processing that word-level factors alone cannot fully account for. In this study, we investigate whether the attention mechanism of Transformer Grammar (TG), which uniquely operates on syntactic structures as representational units, can serve as a cognitive model of human memory retrieval, using Normalized Attention Entropy (NAE) as a linking hypothesis between model behavior and human processing difficulty. Our experiments demonstrate that TG’s attention achieves superior predictive power for self-paced reading times compared to vanilla Transformer’s, with further analyses revealing independent contributions from both models. These findings suggest that human sentence processing involves dual memory representations—one based on syntactic structures and another on token sequences—with attention serving as the general retrieval algorithm, while highlighting the importance of incorporating syntactic structures as representational units.
@@
Ryo Yoshida Shinnosuke Isono Kohei Kajikawa Taiga Someya Yushi Sugimito Yohei Oseki The University of Tokyo Osaka University {yoshiryo0617,isono-shinnosuke,kohei-kajikawa,taiga98-0809,oseki}@g.ecc.u-tokyo.ac.jp sugimoto.yushi.hmt@osaka-u.ac.jp
@@ 1 Introduction
Whether language models (LMs) developed in natural language processing (NLP) are plausible as cognitive models of human sentence processing is a central question in computational psycholinguistics. Over the past two decades, this question has been primarily addressed from the perspective of expectation-based theories—one of the two major classes of human sentence processing theory—examining whether LMs’ next-word prediction can serve as a model of human predictive processing (Hale, 2001; Levy, 2008; Wilcox et al., 2020; Merkx and Frank, 2021; inter alia).
@@
The recent success of Transformer architectures Vaswani et al. (2017) in NLP has unexpectedly opened a new avenue of investigation from the perspective of memory-based theories, the other major class of sentence processing theory. Researchers have proposed that the attention mechanism, despite its engineering origins, can implement a human memory retrieval theory known as cue-based retrieval Van Dyke and Lewis (2003). Recent studies have revealed intriguing parallels between the weighted reference patterns exhibited by the attention mechanism and the elements that humans may retrieve during online sentence comprehension (Ryu and Lewis, 2021; Oh and Schuler, 2022; Timkey and Linzen, 2023).
@@
Computational psycholinguistics has also established that human sentence processing cannot be fully explained by word-level factors alone; rather, syntactic structures have provided compelling explanations for it. For instance, next-word prediction from LMs that explicitly incorporate syntactic structure building demonstrates superior performance in accounting for human brain activity compared to vanilla RNNs and Transformers (Hale et al., 2018; Wolfman et al., 2024); the number of syntactic nodes hypothesized to be constructed per word correlates significantly with both reading times (Kajikawa et al., 2024) and neural activity patterns (Brennan et al., 2012).
@@
Given these findings, if attention can serve as a general algorithm for memory retrieval in human sentence processing, human memory retrieval should be captured by the attention mechanism operating on syntactic structures as well as that operating on token sequences. In this study, we investigate whether the attention mechanism of Transformer Grammar (TG; Sartran et al., 2022), which uniquely operates on syntactic structures as representational units, can serve as a cognitive model of human memory retrieval, using Normalized Attention Entropy (NAE; Oh and Schuler, 2022) as the linking hypothesis between model behavior and human processing difficulty. Our experiments demonstrate that TG’s attention achieves superior predictive power for self-paced reading times compared to vanilla Transformer’s, with further analyses revealing independent contributions from both models. These findings suggest that human sentence processing involves dual memory representations—one based on syntactic structures and another on token sequences—with attention serving as the general retrieval algorithm, while highlighting the importance of incorporating syntactic structures as representational units.
@@ 2.1 Normalized Attention Entropy (NAE)
In recent computational psycholinguistics, attempts have been made to interpret the attention mechanism—a weighted reference of preceding tokens based on Query and Key vectors—as a computational implementation of cue-based retrieval. Notably, Ryu and Lewis (2021) proposed Attention Entropy (AE) as a linking hypothesis, where the diffuseness of attention weights is assumed to quantify the degree of retrieval interference. While AE was initially proposed for modeling interference effects in specific constructions, Oh and Schuler (2022) extended it to naturally occurring text by introducing two normalizations: (i) division by the maximum entropy achievable given the number of preceding tokens, and (ii) sum-to-1 renormalization of attention weights over 1–-th tokens (Normalized AE, NAE).111Oh and Schuler (2022) showed that regression models for predicting reading times fail to converge with vanilla AE.
where represents the attention weight vector when using the -th token as the query in the -th head of layer .222Oh and Schuler (2022) explored NAE calculation using various attention weight formulations, but in this study, we adopt the norm-based attention weight formulation (Kobayashi et al., 2020), which achieved the highest predictive power on the self-paced reading time corpus. In this paper, we employ this NAE as a linking hypothesis between attention mechanisms and human memory retrieval.333While Oh and Schuler (2022) also proposed other metrics based on distances between attention weights at consecutive time steps, we exclusively adopt NAE because (i) in TG, the number of preceding elements varies with time, making distance definition non-trivial, and (ii) Oh and Schuler (2022) demonstrated that NAE’s predictive power subsumes that of distance-based metrics in the self-paced reading time corpus.
@@ 2.2 Transformer Grammar (TG)
Previous research has demonstrated that TG’s probability estimates align more closely with human offline grammaticality judgments (Sartran et al., 2022) and online brain activity than vanilla Transformers (Wolfman et al., 2024). This study investigates whether the attention mechanism of TG, which uniquely operates on syntactic structures as representational units, can serve as a cognitive model of human memory retrieval.
@@ 3.1 NAE calculation with TG
How should the cognitive load of attention from non-lexical tokens (i.e., (X and X)) be attributed to words?
@@
We consider only attention from words, excluding attention from non-lexical tokens.
The adoption of 1-A. is motivated by two factors. First, the self-paced reading time corpus we utilized here provides gold-standard syntactic structures for each sentence, and previous studies have developed predictors based on these annotations (Shain et al., 2020; Isono, 2024). Using the same structural assumptions enables fair comparison with these established predictors, considering the possibility of parsing errors. Second, TG’s current implementation lacks beam search procedure (Stern et al., 2017; Crabbé et al., 2019), an inference technique commonly used in cognitive modeling to handle local ambiguities through parallel parsing (Hale et al., 2018; Sugimoto et al., 2024).444As a proof of concept, we also conducted experiments using multiple syntactic structures generated by word-synchronous beam search with Recurrent Neural Network Grammar (Dyer et al., 2016; Kuncoro et al., 2017; Noji and Oseki, 2021), obtaining similar results (Appendix D).
@@ 3.2 Settings
We used 16-layer, 8-head TG and Transformer (252M parameters).555https://github.com/google-deepmind/transformer_grammars All hyperparameters followed the default settings described in Sartran et al. (2022) (see Appendix A). Following Oh and Schuler (2022), we computed NAE separately for each attention head at the topmost layers and then summed the values across heads.
@@
We used BLLIP-lg, a dataset containing 42M tokens (1.8M sentences) from the Brown Laboratory for Linguistic Information Processing (BLLIP) 1987–89 WSJ Corpus Release 1 (Charniak et al., 2000).666https://catalog.ldc.upenn.edu/LDC2000T43 The corpus was re-parsed using a state-of-the-art constituency parser (Kitaev and Klein, 2018) and split into train-val-test sets by Hu et al. (2020). BLLIP-lg has been widely used for training syntactic LMs, including TG. Following Sartran et al. (2022), we trained a 32K SentencePiece tokenizer (Kudo and Richardson, 2018) on the training set and segmented each sentence into subword units.
Both TG and Transformer were trained at the sentence level: TG maximized the joint probability on action sequences, while Transformer maximized the probability on terminal subword sequences. For training hyperparameters, we largely followed the default settings in Sartran et al. (2022) but adjusted the batch size to fit within the memory constraints of our hardware (NVIDIA A100, 40GB). Accordingly, we tuned other hyperparameters (e.g., learning rate) to maintain training stability. We trained three models with different random seeds and selected the checkpoint with the lowest validation loss for each run.
@@
We used the Natural Stories corpus (Futrell et al., 2018)777https://github.com/languageMIT/naturalstories consisting of 10 stories (485 sentences, 10,245 words) with self-paced reading times collected from 181 anonymized native English speakers. Following Futrell et al.’s preprocessing, data points were removed if (i) a participant scored less than 5/6 on comprehension questions for a story or (ii) individual reading times were less than 100 or greater than 3,000 . Following Oh and Schuler (2022), we also excluded sentence-initial and sentence-final data points. We further removed sentence-second data points, as they lack a log trigram frequency of the previous token required for our baseline regression model. After preprocessing, 724,883 data points from 180 participants remained for statistical analysis, out of the original 848,747 data points.
@@
We evaluated each LM’s NAE contribution to reading time prediction by measuring improvements in regression model fit when adding NAE as predictors. For each model (TG/Transformer), we included both the current word’s NAE (tg_nae/tf_nae) and the previous word’s NAE (tg_nae_so/tf_nae_so) to account for spillover effects.888_so indicates spillover.999Following Oh and Schuler (2022), we summed the subword NAE values for each word. Model improvement was quantified as the increase in log-likelihood (LogLik). This evaluation was conducted for each random seed, and we report the mean LogLik with standard deviation.
The baseline regression model included standard predictors from the Natural Stories corpus:
zone and position: word position in the story and sentence;
wordlen: number of characters in the word;
unigram, bigram, and trigram: log-transformed n-gram frequencies.
@@
tg_surp and tf_surp: surprisal from TG and Transformer;
stack_count: number of elements in the stack (comprising unclosed non-terminals, not-composed terminals, and closed phrases).
Following Oh and Schuler (2022), we included surprisal to test NAE’s significance in the presence of surprisal predictors from the same LMs.101010For an experiment on the predictive power of surprisal itself, see Appendix E. Stack count was included to isolate the cost of holding elements Joshi (1990); Abney and Johnson (1991); Resnik (1992) from their interference effects, which TG’s NAE was designed to capture.
All predictors were z-transformed, and we also included the previous word’s values as predictors to model spillover, except for the positional information. The baseline regression model was a linear mixed-effects model (Baayen et al., 2008) with these fixed effects and by-subjects and by-story random intercepts:
To assess each LM’s independent contribution to reading time prediction, we also conducted likelihood ratio tests by extending Equation 3.2 in two ways: adding both LMs’ NAE versus adding only one LM’s NAE. Note that a larger LogLik from one LM does not necessarily indicate that it contributes above and beyond the other LM, nor does a smaller LogLik indicate no unique contribution. Following Aurnhammer and Frank (2019), we used NAE and surprisal values averaged across random seeds for these nested model comparisons.
@@ 4.1 Does TG’s NAE have predictive power for reading times?
Table 1 presents the contributions of TG’s and Transformer’s NAE to reading time prediction. First, Transformer’s NAE exhibited significant predictive power for reading times, independent of baseline predictors such as surprisal. The effect size was in the expected positive direction (higher NAE values corresponding to longer reading times), primarily showing the immediate effect. This corroborates the arguments of Ryu and Lewis and Oh and Schuler that the attention mechanism—the weighted reference of preceding tokens—functions as a cognitive model of human memory retrieval, despite its engineering-oriented origins.
Second, TG’s NAE exhibited robust predictive power, demonstrating significant positive effects in both immediate and spillover contexts. This finding not only provides additional evidence for incremental construction of syntactic structures in human sentence processing (e.g., Fossum and Levy, 2012), but also suggests that TG’s attention mechanism effectively models memory retrieval from these constructed syntactic representations.
Finally, TG’s NAE made a substantially stronger contribution to reading time prediction (LogLik96.3) compared to Transformer’s NAE (LogLik35.2). This finding suggests that retrieval from syntactic memory representations plays a more dominant role in human sentence processing than retrieval from lexical memory representations. This underscores the importance of incorporating syntactic structures as a unit of memory representation, which we implemented through the integration of TG and NAE here.

### v2

@@ Abstract
Recent work in computational psycholinguistics has revealed intriguing parallels between attention mechanisms and human memory retrieval, focusing primarily on vanilla Transformers that operate on token-level representations. However, computational psycholinguistic research has also established that syntactic structures provide compelling explanations for human sentence processing that token-level factors cannot fully account for. In this paper, we investigate whether the attention mechanism of Transformer Grammar (TG), which uniquely operates on syntactic structures as representational units, can serve as a cognitive model of human memory retrieval, using Normalized Attention Entropy (NAE) as a linking hypothesis between models and humans. Our experiments demonstrate that TG’s attention achieves superior predictive power for self-paced reading times compared to vanilla Transformer’s, with further analyses revealing independent contributions from both models. These findings suggest that human sentence processing involves dual memory representations—one based on syntactic structures and another on token sequences—with attention serving as the general memory retrieval algorithm, while highlighting the importance of incorporating syntactic structures as representational units.
@@
Ryo Yoshida Shinnosuke Isono Kohei Kajikawa Taiga Someya Yushi Sugimoto Yohei Oseki The University of Tokyo NINJAL The University of Osaka {yoshiryo0617,taiga98-0809,oseki}@g.ecc.u-tokyo.ac.jp {s-isono,kohei-kajikawa}@ninjal.ac.jp sugimoto.yushi.hmt@osaka-u.ac.jp
@@ 1 Introduction
Whether language models (LMs) developed in natural language processing (NLP) are plausible as cognitive models of human sentence processing is a central question in computational psycholinguistics. Over the past two decades, this question has been primarily addressed from the perspective of expectation-based theories—one of the two major classes of human sentence processing theory—examining whether LMs’ next-token prediction can serve as a model of human predictive processing (Hale, 2001; Levy, 2008; Wilcox et al., 2020; Merkx and Frank, 2021; inter alia).
@@
The recent success of Transformers Vaswani et al. (2017) in NLP has unexpectedly opened a new avenue of investigation from the perspective of memory-based theories, the other major class of sentence processing theory. Researchers have proposed that the attention mechanism, despite its engineering origins, can implement a human memory retrieval theory known as cue-based retrieval Van Dyke and Lewis (2003). Recent studies have revealed intriguing parallels between the weighted reference patterns exhibited by the attention mechanism and the elements that humans may retrieve during online sentence comprehension (Ryu and Lewis, 2021; Oh and Schuler, 2022; Timkey and Linzen, 2023).
@@
Computational psycholinguistics has also established that human sentence processing cannot be fully explained by token-level factors; rather, syntactic structures have provided compelling explanations for it. For instance, next-token prediction from LMs that explicitly incorporate syntactic structure building demonstrates superior performance in accounting for human brain activity compared to vanilla RNNs and Transformers (Hale et al., 2018; Wolfman et al., 2024); the number of syntactic nodes hypothesized to be constructed per word correlates significantly with both reading times and neural activity patterns (Kajikawa et al., 2024; Brennan et al., 2012).
@@
Given these findings, if attention can serve as a general algorithm for memory retrieval in human sentence processing, human memory retrieval should be captured by the attention mechanism operating on syntactic structures as well as that operating on token sequences. In this paper, we investigate whether the attention mechanism of Transformer Grammar (TG; Sartran et al., 2022), which uniquely operates on syntactic structures as representational units, can serve as a cognitive model of human memory retrieval, using Normalized Attention Entropy (NAE; Oh and Schuler, 2022) as the linking hypothesis between models and humans. Our experiments demonstrate that TG’s attention achieves superior predictive power for self-paced reading times compared to vanilla Transformer’s, with further analyses revealing independent contributions from both models. These findings suggest that human sentence processing involves dual memory representations—one based on syntactic structures and another on token sequences—with attention serving as the general memory retrieval algorithm, while highlighting the importance of incorporating syntactic structures as representational units.111Code for reproducing our results is available at https://github.com/osekilab/TG-NAE.
@@ 2.1 Cue-based retrieval
2.2 Normalized Attention Entropy (NAE)
In recent computational psycholinguistics, attempts have been made to interpret the attention mechanism—a weighted reference of preceding tokens based on Query and Key vectors—as a computational implementation of cue-based retrieval. Notably, Ryu and Lewis (2021) proposed Attention Entropy (AE) as a linking hypothesis, where the diffuseness of attention weights is assumed to quantify the degree of retrieval interference. While AE was initially proposed for modeling interference effects in specific constructions, Oh and Schuler (2022) extended it to naturally occurring text by introducing two normalizations: (i) division by the maximum entropy achievable given the number of preceding tokens, and (ii) sum-to-1 renormalization of attention weights over preceding tokens (Normalized AE, NAE).222Oh and Schuler (2022) showed that regression models for predicting reading times fail to converge with vanilla AE.
where is the set of preceding token positions,333For vanilla Transformers, . is the renormalized attention weight, and represents the attention weight from the -th token (Query) to the -th preceding token (Key) in the -th head of layer .444Oh and Schuler (2022) explored NAE calculation using various attention weight formulations, but in this paper, we adopt the norm-based attention weight formulation (Kobayashi et al., 2020), which achieved the highest predictive power on the self-paced reading time corpus. In this paper, we employ this NAE as a linking hypothesis between attention mechanisms and human memory retrieval.555While Oh and Schuler (2022) also proposed other metrics based on distances between attention weights at consecutive time steps, we exclusively adopt NAE because (i) in TG, the number of preceding elements varies with time, making distance definition non-trivial, and (ii) Oh and Schuler (2022) demonstrated that NAE’s predictive power subsumes that of distance-based metrics in the self-paced reading time corpus.
@@ 2.3 Transformer Grammar (TG)
Previous research has demonstrated that TG’s probability estimates align more closely with human offline grammaticality judgments (Sartran et al., 2022) and online brain activity (Wolfman et al., 2024) than vanilla Transformers. This paper investigates whether the attention mechanism of TG, which uniquely operates on syntactic structures as representational units, can serve as a cognitive model of human memory retrieval.
@@ 3.1 NAE calculation with TG
How should the cognitive load of attention from non-lexical symbols (i.e., (X and X)) be attributed to lexical tokens?
@@
We consider only attention from lexical tokens, excluding attention from non-lexical symbols.
The adoption of 1-A. is motivated by two factors. First, the self-paced reading time corpus we utilized here provides gold-standard syntactic structures for each sentence, and previous studies have developed predictors based on these annotations (Shain et al., 2020; Isono, 2024). Using the same structural assumptions enables fair comparison with these established predictors, considering the possibility of parsing errors. Second, TG’s current implementation lacks beam search procedure (Stern et al., 2017; Crabbé et al., 2019), an inference technique commonly used in cognitive modeling to handle local ambiguities through parallel parsing (Hale et al., 2018; Sugimoto et al., 2024).666As a proof of concept, we also conducted experiments using multiple syntactic structures generated by word-synchronous beam search with Recurrent Neural Network Grammar (Dyer et al., 2016; Kuncoro et al., 2017; Noji and Oseki, 2021), obtaining similar results (Appendix D).
@@ 3.2 Settings
We used 16-layer, 8-head TG and Transformer (252M parameters).777https://github.com/google-deepmind/transformer_grammars All hyperparameters followed the default settings described in Sartran et al. (2022) (see Appendix A). Following Oh and Schuler (2022), we computed NAE separately for each attention head at the top layer and then summed the values across heads.
@@
We used BLLIP-lg, a dataset containing 42M tokens (1.8M sentences) from the Brown Laboratory for Linguistic Information Processing (BLLIP) 1987–89 WSJ Corpus Release 1 (Charniak et al., 2000).888https://catalog.ldc.upenn.edu/LDC2000T43 The corpus was re-parsed using a state-of-the-art constituency parser (Kitaev and Klein, 2018) and split into train-val-test sets by Hu et al. (2020). BLLIP-lg has been widely used for training syntactic LMs, including TG. Following Sartran et al. (2022), we trained a 32K SentencePiece tokenizer (Kudo and Richardson, 2018) on the training set and segmented each sentence into subword units.
Both TG and Transformer were trained at the sentence level: TG maximized the joint probability on action sequences, while Transformer maximized the probability on terminal subword sequences. For training hyperparameters, we largely followed the default settings in Sartran et al. (2022) but adjusted the batch size to fit within the memory constraints of our hardware (NVIDIA A100, 40GB). Accordingly, we tuned other hyperparameters (e.g., learning rate) to maintain training stability.999For the detailed hyperparameters, see Appendix A. We trained three models with different random initialization seeds and selected the checkpoint with the lowest validation loss for each run.
@@
We used the Natural Stories corpus (Futrell et al., 2018),101010https://github.com/languageMIT/naturalstories. We used the corrected version that addresses the data misalignment issue identified in May 2025. “a series of English narrative texts designed to contain many low-frequency and psycholinguistically interesting syntactic constructions while still sounding fluent and coherent.” We selected this corpus for these “interesting” syntactic constructions, which provide an ideal testbed for investigating memory retrieval effects that might be less pronounced in simpler, more naturalistic texts. The corpus has also been used in several studies investigating memory-related processing mechanisms (Shain et al., 2016; Dotlačil, 2021; Isono, 2024).
The Natural Stories corpus consists of 10 stories (485 sentences, 10,245 words) with self-paced reading times collected from 181 anonymized native English speakers. Following Futrell et al.’s preprocessing, data points were removed if (i) a participant scored less than 5/6 on comprehension questions for a story or (ii) individual reading times were less than 100 or greater than 3,000 . Following Oh and Schuler (2022), we also excluded sentence-initial and sentence-final data points. We further removed sentence-second data points, as they lack the log trigram frequency of the previous token required for our baseline regression model. After preprocessing, 725,875 data points from 180 participants remained for statistical analysis, out of the original 848,875 data points.
@@
We evaluated each LM’s NAE contribution to reading time prediction by measuring improvements in regression model fit when adding NAE as predictors. For each model (TG/Transformer), we included both the current word’s NAE (tg_nae/tf_nae) and the previous word’s NAE (tg_nae_so/tf_nae_so) to account for spillover effects (Mitchell, 1984).111111_so indicates spillover.121212Following Oh and Schuler (2022), we summed the subword NAE values for each word. Model improvement was quantified as the increase in log-likelihood (LogLik). This evaluation was conducted for each random seed, and we report the mean LogLik with standard deviation.
Following previous studies such as Dotlačil (2021), Shain et al. (2016), and Isono (2024), the baseline regression model controlled for non-structural, basic aspects of text known to affect reading times:
zone and position (integer): word position in the story and sentence;
wordlen (integer): number of characters in the word;
unigram, bigram, and trigram (continuous): log-transformed n-gram frequencies.
@@
tg_surp and tf_surp (continuous): surprisal from TG and Transformer;
stack_count (integer): number of elements in the stack (comprising unclosed non-terminals, not-composed terminals, and closed phrases).
Following Oh and Schuler (2022), we included surprisal to test NAE’s significance in the presence of surprisal predictors from the same LMs.131313For an experiment on the predictive power of surprisal itself, see Appendix E. Stack count was included to isolate the cost of holding elements Joshi (1990); Abney and Johnson (1991); Resnik (1992) from their interference effects, which TG’s NAE was designed to capture. For the correlations between the predictors, see Appendix B.
All predictors were z-transformed, and we also included the previous word’s values as predictors to model spillover, except for the positional information. The baseline regression model was a linear mixed-effects model (Baayen et al., 2008) with these fixed effects and by-subject and by-story random intercepts:
To assess each LM’s independent contribution to reading time prediction, we also conducted likelihood ratio tests (Wurm and Fisicaro, 2014) by extending Equation 3.2 in two ways: adding both LMs’ NAE versus adding only one LM’s NAE. Note that a larger LogLik from one LM does not necessarily indicate that it contributes above and beyond the other LM, nor does a smaller LogLik indicate no unique contribution. Following Aurnhammer and Frank (2019), we used NAE and surprisal values averaged across random seeds for these nested model comparisons.
@@ 4.1 Does TG’s NAE have predictive power for reading times?
Table 1 presents the contributions of TG’s and Transformer’s NAE to reading time prediction. First, Transformer’s NAE exhibited significant predictive power for reading times, independent of baseline predictors such as surprisal. The effect size was in the expected positive direction (higher NAE values corresponding to longer reading times), showing both immediate and spillover effects. This corroborates the arguments of Ryu and Lewis (2021) and Oh and Schuler (2022) that the attention mechanism—the weighted reference of preceding tokens—functions as a cognitive model of human memory retrieval, despite its engineering-oriented origins.
Second, TG’s NAE exhibited robust predictive power, demonstrating significant positive effects in both immediate and spillover conditions. This finding not only provides additional evidence for incremental construction of syntactic structures in human sentence processing (e.g., Fossum and Levy, 2012), but also suggests that TG’s attention mechanism effectively models memory retrieval from these constructed syntactic representations.
Finally, TG’s NAE made a substantially stronger contribution to reading time prediction (LogLik76.6) compared to Transformer’s NAE (LogLik42.8). This finding suggests that retrieval from syntactic memory representations plays a more dominant role in human sentence processing than retrieval from lexical memory representations. This underscores the importance of incorporating syntactic structures as a unit of memory representation, which we implemented through the integration of TG and NAE here.